# Operator Deep Smoothing for Implied Volatility

### v1

@@ Appendix A Neural Operators
We give a review of the kernel integral transform neural operator framework of [20], and expand in more detail on its graph neural operator.
@@
Let $\mathcal{A}$ and $\mathcal{U}$ be input and output space of an operator learning problem, as introduced in Section 3.1. Then, $\mathcal{A}$ and $\mathcal{U}$ are Banach spaces of functions $D\to\mathbb{R}^{c_{\text{in}}}$ and $D\to\mathbb{R}^{c_{\text{out}}}$, respectively, where $D$ is a (bounded) domain in $\mathbb{R}^{d}$. The mathematical analysis of neural operators in [20] summarized hereafter, as well as the definition of universality and discretization-invariance in Section 3.1, makes use of the following terms and notations.
In the context of [20], a domain is a bounded and connected open set that is topologically regular (in the sense that it is the interior of its closure). A domain $D$ is Lipschitz if its boundary locally is the graph of a Lipschitz continuous function defined on an open ball of $\mathbb{R}^{d-1}$. An open ball – for any metric space $\mathcal{X}=(\mathcal{X},d)$ – is the set
@@
It is well-known that this topology on $C(\mathcal{A},\mathcal{U})$ is induced by the metric
@@
Therefore, the notion of density in $C(\mathcal{A},\mathcal{U})$, as used to define universality of neural operators in Section 3.1, is well-defined.
@@ A.1 Kernel Integral Neural Operators
    • The bias term $b$ is itself a function from $D$ to $\mathbb{R}^{c_{j+1}}$,
@@
Theorem A.1 (Universal Approximation; Theorem 11 of [20]).
@@
$\tilde{L}(h|_{\pi})(y)=\sigma\left(Wv(y)+\sum_{l=1}^{p}\kappa(y,x_{l})h(x_{l})\lambda_{d}(D_{l})+b(y)\right),\quad y\in D.$ (13)
[20] establishes the following.
Theorem A.2 (Discretization Invariance; Theorem 8 of [20]).
@@
A.2 Graph Neural Operator
The curse of dimensionality makes the direct implementation (14) prohibitively expensive in practice. Instead, [3] introduces graph neural operators (or, GNOs, for short) which replace the kernel integral operation at the heart of the framework by a sum approximation and organizes the constituent terms using a directed graph structure: The discretization $\pi=\{x_{1},\dots,x_{m}\}$ of the input data $\mathbf{a}$ is enriched with a directed graph structure $G_{\mathbf{a}}=(V,E)$, allowing the following implementation $\tilde{F}^{\theta}$ of $F^{\theta}$:
$\tilde{L}(v|_{\pi})(y)=\sigma\left(Wv(y)+\frac{1}{|\mathcal{N}_{\text{in}}(y)|}\sum_{x\in\mathcal{N}_{\text{in}}(y)}\kappa(y,x)v(x)+b(y)\right),\quad y\in V.$ (15)
Here, $\mathcal{N}_{\text{in}}(y)$ is the set of so-called in-neighbors of $y$ in the graph $G_{\mathbf{a}}$: $x\in\mathcal{N}_{\text{in}}(y)$ iff $(x,y)\in E$. It is clear that, to compute output at $y$, the point $y$ must be included as a node into the graph $G_{\mathbf{a}}$. It is important to reconcile the input and output locations of the layers when creating the graph structure to enable an efficient implementation using message passing algorithms.
@@ C.1 Data
Our numerical experiments are based on the "Option Quotes" dataset product available for purchase from the CBOE. Our version of the dataset contains relevant data features for S&P 500 Index options for the years 2012 through 2021 and is summarized on a 20-minute interval basis.
@@
We compute simple mids for options and underlying by averaging bid and ask quotes and use this aggregate as a reference price for all our subsequent computations. We calculate discount factors and forward prices from Put-Call parity using the industry standard technique based on linear regression. We compute time-to-expiry in units of one year as well as log-moneyness as defined in Section 2. We extract implied volatilities using the py-vollib-vectorized project available in the Python Package Index at the location https://pypi.org/project/py-vollib-vectorized/. py-vollib-vectorized implements a vectorized version of [19]’s Let’s-be-rational state-of-the-art method for computing implied volatility. We discard all implied volatilities of in-the-money options, or, in other words, we compose our implied volatility surface from Put options for non-positive log-moneyness values and Call options for positive log-moneyness values.
@@ C.3 Training
where $\mathcal{V}(x,v(x))$ is the Black-Scholes Vega defined in Section 1 (adjusted for transformed coordinates).
@@
The particular weighting of the individual terms has been retrieved by manual experimentation, led by the findings of [2].
@@ C.4 Evaluation
Here we provide additional results to supplement our performance evaluation of operator deep smoothing (OpDS) for implied volatility. Table 2 and Table 3 display descriptive statistics of OpDS versus SVI. We also included a comparison to a less sophisticated version of OpDS without monthly finetuning, which we denote by OpDS*. Moreover, we include Figures 5 and 7, which are versions of Figures 3 and  4 with the only difference being that Figures 5 and 7 were created using OpDS* instead of OpDS. Finally, Figure 6 shows the average spatial distribution of benchmark metrics and arbitrage term over the training dataset, which complements the same averages on the test dataset shown in Figures 3.
@@ D.1 Example Plots
We plot the results of operator deep smoothing (OpDS) vs. SVI on example inputs. To aid presentation, we include only every thrid market datapoint into the plots.
Figure 8: Smoothing of quotes $\mathbf{v}$ from 20.07.2012 at 10:50:00.
Figure 9: Smoothing of quotes $\mathbf{v}$ from 21.10.2016 at 13:10:00.
Figure 10: Smoothing of quotes $\mathbf{v}$ from 04.01.2021 at 10:50:00.
@@ Appendix E SVI
SVI, originally devised by Gatheral [10], stands for Stochastic Volatility Inspired and is a low-dimensional parametrization for implied volatility slices (namely for each maturity). It captures the key features of implied volatility of Equity indices, and has become the industry-wide benchmark for implied volatility smoothing on such markets. Its "raw" variant parametrizes the "slice" of implied volatility at a given time-to-expiry $\tau$ as follows:
@@
While stylistically accurate, SVI does not easily guarantee absence of static arbitrage opportunities, and several authors have investigated this issue [11, 27, 29, 28]. To produce our SVI benchmark we therefore rely on the constrained SLSQP optimizer provided by the SciPy scientific computing package for Python, with the mean square error objective, a positivity constraint and the constraint (3) (computed in closed form), and the following parameter bounds:

### v2

@@ Appendix A Neural Operators
We give a review of the kernel integral transform neural operator framework of Kovachki et al., (2023), and expand in more detail on its graph neural operator.
@@
Let $\mathcal{A}$ and $\mathcal{U}$ be input and output space of an operator learning problem, as introduced in Section 3.1. Then, $\mathcal{A}$ and $\mathcal{U}$ are Banach spaces of functions $D\to\mathbb{R}^{c_{\text{in}}}$ and $D\to\mathbb{R}^{c_{\text{out}}}$, respectively, where $D$ is a (bounded) domain in $\mathbb{R}^{d}$. The mathematical analysis of neural operators in Kovachki et al., (2023) summarized hereafter, as well as the definition of universality and discretization-invariance in Section 3.1, make use of the following terms and notations.
In the context of Kovachki et al., (2023), a domain is a bounded and connected open set that is topologically regular (in the sense that it is the interior of its closure). A domain $D$ is Lipschitz if its boundary locally is the graph of a Lipschitz continuous function defined on an open ball of $\mathbb{R}^{d-1}$. An open ball – for any metric space $\mathcal{X}=(\mathcal{X},d)$ – is the set
@@
It is well known that this topology on $C(\mathcal{A},\mathcal{U})$ is induced by the metric
@@
Therefore, the notion of density in $C(\mathcal{A},\mathcal{U})$, as used to define universality of neural operators in Section 3.1, is well defined.
@@ A.1 Kernel Integral Neural Operators
    • The bias term $b_{j}$ is itself a function from $D$ to $\mathbb{R}^{c_{j+1}}$,
@@
Theorem A.1 (Universal Approximation; Theorem 11 of Kovachki et al., (2023)).
@@
$\tilde{L}(h|_{\pi})(y)=\sigma\left(Wh(y)+\sum_{l=1}^{p}\kappa(y,x_{l})h(x_{l})\lambda_{d}(D_{l})+b(y)\right),\quad y\in D.$ (12)
Kovachki et al., (2023) establishes the following.
Theorem A.2 (Discretization Invariance; Theorem 8 of Kovachki et al., (2023)).
@@
A.2 Graph Neural Operators
The curse of dimensionality makes the direct implementation (13) prohibitively expensive in practice. Instead, Anandkumar et al., (2020) introduces graph neural operators (or, GNOs, for short) which replace the kernel integral operation at the heart of the framework by a sum approximation and organizes the constituent terms using a directed graph structure: The discretization $\pi=\{x_{1},\dots,x_{m}\}$ of the input data $\mathbf{h}=h|_{\pi}$ is enriched with a directed graph structure $G_{\mathbf{h}}=(V,E)$, allowing the following implementation $\tilde{F}^{\theta}$ of $F^{\theta}$:
$\tilde{L}(h|_{\pi})(y)=\sigma\left(Wh(y)+\frac{1}{|\mathcal{N}_{\text{in}}(y)|}\sum_{x\in\mathcal{N}_{\text{in}}(y)}\kappa(y,x)h(x)+b(y)\right),\quad y\in V.$ (14)
Here, $\mathcal{N}_{\text{in}}(y)$ is the set of so-called in-neighbors of $y$ in the graph $G_{\mathbf{h}}$: $x\in\mathcal{N}_{\text{in}}(y)$ iff $(x,y)\in E$. It is clear that, to compute output at $y$, the point $y$ must be included as a node into the graph $G_{\mathbf{h}}$. On the other hand, it is necessary to drop the local linear transformation with $W$ if $y$ is not part of the input data locations $\pi$ (compare our discussion in 3). It is important to reconcile the input and output locations of the layers when creating the graph structure to enable an efficient implementation using message passing algorithms.
@@ C.1 Data
Our numerical experiments are based on the ”Option Quotes” dataset product available for purchase from the CBOE. Our version of the dataset contains relevant data features for S&P 500 Index options for the years 2012 through 2021 and is summarized on a 20-minute interval basis.
@@
We compute simple mids for options and underlying by averaging bid and ask quotes and use this aggregate as a reference price for all our subsequent computations. We calculate discount factors and forward prices from Put-Call parity using the industry standard technique based on linear regression. We compute time-to-expiry in units of one year as well as log-moneyness as defined in Section 2. We extract implied volatilities using the py-vollib-vectorized project available in the Python Package Index at the location https://pypi.org/project/py-vollib-vectorized/. py-vollib-vectorized implements a vectorized version of Jäckel, (2015)’s Let’s-be-rational state-of-the-art method for computing implied volatility. We discard all implied volatilities of in-the-money options, or, in other words, we compose our implied volatility surface from Put options for non-positive log-moneyness values and Call options for positive log-moneyness values.
@@ C.3 Training
where $\mathcal{V}(x,v(x))$ is the Black-Scholes Vega, the sensitivity of the Black-Scholes option price with respect to its volatility parameter:
defined in Section 1 (adjusted for transformed coordinates).
@@
The particular weighting of the individual terms has been retrieved by manual experimentation, led by the findings of Ackerer et al., (2020).
@@ C.4 Evaluation
Here we provide additional results to supplement our performance evaluation of operator deep smoothing (OpDS) for implied volatility. Table 3 and Table 4 display descriptive statistics of OpDS versus SVI. OpDS* refers to benchmarking without monthly finetuning. Moreover, we include Figures 5 and 7, which are versions of Figures 3 and  4 with the only difference being that Figures 5 and 7 were created using OpDS* instead of OpDS. Finally, Figure 6 shows the average spatial distribution of benchmark metrics and arbitrage term over the training dataset, which complements the same averages on the test dataset shown in Figures 3.
@@
C.5 Comparison to Classical Neural Networks
We reproduce the experiment underlying Table 1 of Ackerer et al., (2020) using our operator deep smoothing approach. Given observed option quotes, it involves dropping 50% of datapoints, and then measuring the MAPE of the smoothed surface at retained datapoints (“Train”) as well as dropped datapoints (”Test”). “Interpolate” and “Extrapolate” are different settings dictating how exactly the datapoints which to drop are selected (for details refer to Ackerer et al., (2020)). The experiment is performed on end-of-day S&P 500 data in the period from January to April 2018 and averaged percentiles are reported. Table 5 reproduces the relevant row from Table 1 of Ackerer et al., (2020) (“DS”) as well as our results averaged over 25 repetitions (“OpDS”).
Table 5: Backtesting results of Operator Deep Smoothing vs. Deep Smoothing (“DS”; taken from [2]); quantiles in %, Jan-Apr 2018 end-of-day SPX data.
Interpolation Extrapolation
MAPE Train Test Train Test
$\lambda$ $q_{05}$ $q_{50}$ $q_{95}$ $q_{05}$ $q_{50}$ $q_{95}$ $q_{05}$ $q_{50}$ $q_{95}$ $q_{05}$ $q_{50}$ $q_{95}$
OpDS 10 0.5 0.7 1.0 0.5 0.7 1.1 0.5 0.7 1.0 0.7 0.9 1.3
DS 10 0.5 0.7 1.2 0.5 0.8 1.2 0.4 0.6 0.9 1.2 1.7 2.4
@@ D.1 Example Plots
We plot the results of operator deep smoothing (OpDS) vs. SVI on example inputs. To aid the visual clarity of our plots, we display only every third market datapoint.
Figure 8: Smoothing of quotes $\mathbf{v}\in\mathcal{D}_{\text{val}}$ from 20.07.2012 at 10:50:00.
Figure 9: Smoothing of quotes $\mathbf{v}\in\mathcal{D}_{\text{val}}$ from 21.10.2016 at 13:10:00.
Figure 10: Smoothing of quotes $\mathbf{v}\in\mathcal{D}_{\text{train}}$ from 04.01.2021 at 10:50:00.
@@ D.2 Monthly Breakdown of Spatial Distributions of Benchmark Metrics on Test Dataset
D.3 Example Plots: Option Metrics End-Of-Day US Index Options Data
Figure 13: Smoothing of SPX end-of-day data from 07.01.2021. Every third datapoint displayed.
Figure 14: Smoothing of NDX end-of-day data from 07.01.2021. Every second datapoint displayed.
Figure 15: Smoothing of DJX end-of-day data from 07.01.2021. Every datapoint displayed.
Figure 16: Smoothing of RUT end-of-day data from 07.01.2021. Every datapoint displayed.
@@ Appendix E SVI
SVI, originally devised by Gatheral (Gatheral,, 2004), stands for Stochastic Volatility Inspired and is a low-dimensional parametrization for implied volatility slices (namely for each maturity). It captures the key features of implied volatility of Equity indices, and has become the industry-wide benchmark for implied volatility smoothing on such markets. Its ”raw” variant parametrizes the ”slice” of implied volatility at a given time-to-expiry $\tau$ as follows:
@@
While stylistically accurate, SVI does not easily guarantee absence of static arbitrage opportunities, and several authors have investigated this issue (Gatheral and Jacquier,, 2014; Martini and Mingone,, 2022; Mingone,, 2022; Martini and Mingone,, 2023). To produce our SVI benchmark we therefore rely on the constrained SLSQP optimizer provided by the SciPy scientific computing package for Python, with the mean square error objective, a positivity constraint and the constraint (2) (computed in closed form), and the following parameter bounds: